A body of mass 25kg, moving at 3m/s on a rough horizontal floor is brought to rest after sliding through a distance of 2.5m on the floor. What is the coefficient of sliding friction (g=10)


1
Expert's answer
2021-04-15T20:41:00-0400

Answer

Using v=u+at

a = "\frac{3}{t}"

Using s=ut+"0.5at^2"

"d = 0.5a t^2"

"2.5 =0 .5\times\frac{ 3}{t}t^2"

2.5 = 1.5 t

t = 1.67 seconds to stop

then"a = \frac{3}{t }= 1.8 m/s^2"

F = ma

F = 25 "\times" 1.8 = 45 N

weight = 25g = 250N


coef of friction "=\frac{ F}{weight} = \frac{45}{250}

= 0.18"